OpenStudy (anonymous):

PLEASE HELP MEE What is the sum of a 6-term geometric sequence if the first term is 11, the last term is -11,264 and the common ratio is -4?

OpenStudy (anonymous):

Sn = a(1 - r^n) / 1 - r a = first term = 11 n = number of term = 6 r = common ratio = -4 Sn = (11)(1- -4^7) / (1 - -4) =?

OpenStudy (anonymous):

use geometric series formula

OpenStudy (anonymous):

a1=11 r= -4 n=6 S(9)= -9009 @dauspex ?